# Revisiting Reliability in Large-Scale Machine Learning Research Clusters

### v1

@@ VI Related Work
ML Infrastructure. A growing interest in ML has led to work investigating infrastructure failures as well as scheduler effects, specifically for ML clusters [24, 31]. This includes analyzing aspects from job sizes to failure properties, predominantly at maximum job scales reaching into the tens of GPUs. IBM’s AI infrastructure and failure taxonomy was introduced in [18]—we provide a similarly motivated taxonomy and quantify the failures in detail. The design and operation of Meta’s production scheduler, MAST, was also recently studied [15], demonstrating a need for datacenter-level load balancing. The rise of LLMs has recently spurred a number of recent works specifically focusing on reliability for ML training [33, 17, 25], assuming orders of magnitude larger scale than previously seen in general-purpose ML experiments. Our work is unique as the clusters serve both small and large jobs in terms of scale and model architecture diversity and does so in a cluster-level research setting.
Model-System Codesign. LLM-specific parallelization techniques, such as Megatron-LM [39], have been developed—naturally producing a hierarchical mapping of data and model parallelism to common network topologies [44]. LLM failures and mitigations at ByteDance were studied in Megascale [25]. A similar study was performed for Alibaba training systems in Unicron [21] and a specialized HPN network was proposed to mitigate networking-related failures [36]. Llama 3 also demonstrated benefits from software-hardware codesign to improve resilience [33]. In addition, fault-tolerant, resilient training has also been explored for other large deep learning model development [32, 46]. While our work tackles the same model training resilience as well as general purpose cluster reliability, our experience is unique in that we operate at both ends of the scale, motivating solutions that are practical to be deployed in a black-box manner, yet still battle tested at the scale of four thousand GPUs. Finally, our workloads are diverse [41, 43, 28, 16, 34, 38], spanning vision, language, and mixed-modality models in a rapidly evolving research setting.

### v2

@@ VI Related Work
ML Infrastructure. A growing interest in ML has led to work investigating infrastructure failures as well as scheduler effects, specifically for ML clusters [37, 44]. This includes analyzing aspects from job sizes to failure properties, predominantly at maximum job scales reaching into the tens of GPUs. IBM’s AI infrastructure and failure taxonomy was introduced in [29]—we provide a similarly motivated taxonomy and quantify the failures in detail. The design and operation of Meta’s production scheduler, MAST, was also recently studied [20], demonstrating a need for datacenter-level load balancing. The rise of LLMs has recently spurred a number of recent works specifically focusing on reliability for ML training [47, 28, 38], assuming orders of magnitude larger scale than previously seen in general-purpose ML experiments. Recent work has analyzed LLM-specific datacenter workloads [35], finding that evaluation jobs are important. In comparison, our work serves the entire range of job sizes from 1 to 4k GPUs and does so with a variety of cutting-edge machine learning research workloads—emphasizing general purpose reliability techniques.
Model-System Codesign. LLM-specific parallelization techniques, such as Megatron-LM [54], have been developed—naturally producing a hierarchical mapping of data and model parallelism to common network topologies [60]. LLM failures and mitigations at ByteDance were studied in Megascale [38]. A similar study was performed for Alibaba training systems in Unicron [32], and a specialized HPN network was proposed to mitigate networking-related failures [50]. LLaMa 3 also demonstrated benefits from software-hardware codesign to improve resilience [47]. In addition, fault-tolerant, resilient training has also been explored for other large deep learning model development [46, 64]. While our work tackles the same model training resilience as well as general purpose cluster reliability, our experience is unique in that we operate at both ends of the scale, motivating solutions that are practical to be deployed in a black-box manner, yet still battle tested at the scale of four thousand GPUs. Finally, our workloads are diverse [56, 59, 41, 22, 48, 53], spanning vision, language, and mixed-modality models in a rapidly evolving research setting.